I actually have two, but they are similar.
Problem A: Candice and Pat walked to the site in 9 hours. The next morning they jogged home in 3 hours. How far did they jog if they jogged 4 miles per hour faster than they walked?
Problem B: On Monday, Voltaire drove to town at 60 miles per hour. On Tuesday, he drove to town at 40 miles per hour. If the total traveling time for both trips was 15 hours, how far was it to town?
Thanks so much!

Answer by khwang(438)

Sol: (A) If the distance between home and the site is D miles.
The walking speed was D/9 miles/hr and the jogging speed was D/3 miles/hr.
Since the difference of the speed is 4 miles/hr, we have
D/3 = 4 + D/9,
Or 3D = 36 + D, or 2D = 36, so D = 18 miles.
(B) If the daily driving distance to town site is x miles.
On Monday,he spent x/60 hrs and on Monday,he spent x/40 hrs in driving to town.
Since the total travelling time was15 hrs, we get
x/60 + x/40 = 15,
Multiplying by 120 on both side: 2x + 3x = 1800,
Or 5x = 1800, x = 360 miles